OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. cos(pi/3)cos(pi/5)+sin(pi/3)sin(pi/5)

OpenStudy (anonymous):

this is the subtraction angle formula for cosine i.e. it is \[\cos(\frac{\pi}{3}-\frac{\pi}{5})\]

OpenStudy (anonymous):

your job is to compute \[\frac{\pi}{3}-\frac{\pi}{5}\]

jimthompson5910 (jim_thompson5910):

satellite73 is using the identity cos(x-y) = cos(x)*cos(y) + sin(x)*sin(y)

OpenStudy (anonymous):

0.91354545764

OpenStudy (anonymous):

zactly btw don't be confused with the arithmetic finding \[\frac{\pi}{3}-\frac{\pi}{5}\] is identical to \[\frac{1}{3}-\frac{1}{5}\] then stick a \(\pi\) next to it

OpenStudy (anonymous):

skip the decimals work with a fraction

OpenStudy (anonymous):

do you want a single fraction?

OpenStudy (anonymous):

yes, a single fraction although it is not really what I want, it is what the question wants personally i don't care too much one way or the other but in any case find \(\frac{1}{3}-\frac{1}{5}\) put a \(\pi\) next to it then write \[\cos(\text{whatever})\]

OpenStudy (anonymous):

let me know what you get and i will check it

OpenStudy (anonymous):

2/15pi?

OpenStudy (anonymous):

right so final answer is \[\cos\left(\frac{2\pi}{15}\right)\]

OpenStudy (anonymous):

Thank you so very much:)

OpenStudy (anonymous):

yw